OpenStudy (anonymous):

Name the opposite and reciprocal for. 1. N/2 2. 1.3

OpenStudy (anonymous):

1.)=N/2=2/N for reciprocal 2.)1.3=1+3/10=13/10=10/13 is the reciprocal

OpenStudy (anonymous):

and the opposite i believe its the negative of 1 and 2 which is -N/2 and -1.3